OpenStudy (anonymous):

Eve bought a book for $10 and then spent half of her remaining money on a train ticket. She then bought lunch for $4 and spent half of her remaining money at a bazaar. She left the bazaar with $20. How much did she start with?

OpenStudy (hoblos):

before the bazaar she had 20*2 = $40 before buying the lunch she had 40 +4 = $44 before buying the train ticket she had 44*2 = $88 before buying the book she had 88+10 = $98 she started with $98

OpenStudy (anonymous):

Thank you! :)